Isopropyl alcohol is mixed with water to produce a solution that is 31.0% alcohol by volume.
How many milliliters of each component are present in 635 mL of this solution?
1
Expert's answer
2019-09-27T04:45:42-0400

"V_{alc}=\frac{\psi_{alc} V{sol}}{100}=\frac{31.0*635}{100}=" 196.85mL

"V_{aq}=V_{sol}-V_{alc}=635-196.85=438.15mL"

Answer "V_{aq}=438.15mL;V_{alc}=196.85mL"